# On the asymptotic profile of solutions to semilinear damped wave equations with critical nonlinearities

### v1

@@ 2.1. Philosophy of our approach
$\left|\mathcal{N}(u(\tau,\cdot))-\mathcal{N}(v(\tau,\cdot))\right|$ $=\left|(u-v)(\tau,\cdot)\times\left(\int_{0}^{1}\left(\frac{\partial\mathcal{N}}{\partial s}\right)(v+\kappa(u-v))d\kappa\right)(\tau,\cdot)\right|$
$\lesssim|(u-v)(\tau,\cdot)|\times\left|\left(\int_{0}^{1}|v+\kappa(u-v)|^{\frac{2}{n}}\mu\left(|v+\kappa(u-v)|\right)d\kappa\right)(\tau,\cdot)\right|$
$\lesssim|(u-v)(\tau,\cdot)|\left(|u(\tau,\cdot)|^{\frac{2}{n}}+|v(\tau,\cdot)|^{\frac{2}{n}}\right)\int_{0}^{1}\|\mu(v+\kappa(u-v))(\tau,\cdot)\|_{L^{\infty}}d\kappa.$
@@
$\mathcal{J}_{1}(t)$ $:=\int_{0}^{t/2}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}(1+\tau)^{-1+\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau,$
$\mathcal{J}_{2}(t)$ $:=\int_{t/2}^{t}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}(1+\tau)^{-1+\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau.$
We can easily see that
$\mathcal{J}_{1}(t)$ $\lesssim(1+t)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\int_{0}^{t/2}(1+\tau)^{-1+\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\lesssim\int_{0}^{t/2}(1+\tau)^{-1}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau\lesssim\mathcal{I}(\varepsilon_{0}),$
and
$\mathcal{J}_{2}(t)$ $\lesssim\int_{t/2}^{t}(1+t-\tau)^{-1}\mu\left(\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}\right)d\tau\lesssim\mathcal{I}(\varepsilon_{0}),$
for all $t>0$ due to $n\left(1/m-1/2\right)<2$ and $1+\tau\geq 1+t-\tau$ for $\tau\in[t/2,t]$. Thus, we completed the proof of Proposition 2.3. ∎
Proof of Theorem 1.1. Using again Lemma 2.1, it is the fact that
@@
for all $T>0$. For this reason, combining with Propositions 2.2 and 2.3, we obtain the following two important estimates for all $u,v\in X(T,\varepsilon_{0}),\,\varepsilon_{0}\in(0,1]$:
@@
for all $\varepsilon\in(0,\bar{\varepsilon}]$. Moreover, using (20) implies that $\{u_{j}\}_{j=0}^{\infty}$ is a Cauchy sequence in the Banach space $Y(T)$. Furthermore, $X(T,\varepsilon_{0})$ is a closed subset of $Y(T)$ with respect to the metric $Y(T)$ (see Lemma 2.2). Therefore, there exists a solution $u=\Phi[u]$ in $X(T,\varepsilon_{0})$ for all $T>0$. Since $T$ is arbitrary, the solution is global, that is, $u\in X(\infty,\varepsilon_{0})$.
Next, we need to show that $u\in\mathcal{C}([0,\infty),H^{2}\cap L^{\alpha}\cap L^{\infty})$. To prove this property, we recall that
$u(t,x)=\Phi[u](t,x)$ $=u^{\rm lin}(t,x)+\int_{0}^{t}\mathcal{K}(t-\tau,x)\ast_{x}\mathcal{N}(u(\tau,x))d\tau.$
Since the linear part $u^{\rm lin}(t,x)$ obviously satisfies continuity, it suffices to show that
@@
Finally, we establish the uniqueness of the global solution $u$ , which lies in the space $\mathcal{C}([0,\infty),H^{2}\cap L^{\alpha}\cap L^{\infty})$. Let $u,v$ be solutions to (1) belonging to this space. Using again the arbitrary positive number $T$ and performing the same proof steps of Proposition 2.3, we obtain
$\|u-v\|_{Y(t)}\lesssim\int_{0}^{t}\|u-v\|_{Y(\tau)}\big(\|u\|_{X(\tau)}^{\frac{2}{n}}+\|v\|_{X(\tau)}^{\frac{2}{n}}\big)d\tau,$
for all $t\in[0,T]$. One can easily see that there exists a constant $M(T)$ such that $\|u\|_{X(T)}^{\frac{2}{n}}+\|v\|_{X(T)}^{\frac{2}{n}}\leq M(T)$. Therefore, we immediately have
@@ 3. Asymptotic behaviors of global solution
Let $1\leq\rho\leq q<\infty$, $q\neq 1$, $\beta_{q}:=(n-1)\big|\frac{1}{2}-\frac{1}{q}\big|$ and $s_{1}\geq s_{2}\geq 0$. Then, the following estimate holds for $t\geq 1$ (see Theorem 1.2 in [14]):
@@
$\hskip 56.9055pt\lesssim(1+t)^{-\frac{n}{2}(\frac{1}{\rho}-\frac{1}{q})-\frac{s_{1}-s_{2}}{2}-1}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}}+e^{-ct}\||\nabla|^{s_{1}}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{q}-1}_{q}},$
@@
$\hskip 56.9055pt\lesssim(1+t)^{-\frac{n}{2\rho}-\frac{s_{1}-s_{2}}{2}-1}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}}+e^{-ct}\||\nabla|^{s_{1}+d}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{r}-1}_{r}}.$
@@
We need to prove the estimate of the $L^{\infty}$ norm. Proposition 2.10 in [14] implies
$\||\nabla|^{s_{1}}\big(\chi_{\rm L}(|\nabla|)\mathcal{K}(t,x)-\mathcal{G}(t,x)\big)\ast_{x}\varphi(x)\|_{L^{\infty}}\lesssim(1+t)^{-\frac{n}{2\rho}-\frac{s_{1}-s_{2}}{2}-1}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}},$
for all $1\leq\rho\leq\infty$ and $s_{1}\geq s_{2}\geq 0$. In addition, from Proposition 2.5 in [14], we obtain
Moreover, it is the fact that
$\quad\lesssim\||\nabla|^{s_{1}}\langle\nabla\rangle^{d}\chi_{\rm H}(|\nabla|)\mathcal{K}(t,x)\ast_{x}\varphi(x)\|_{L^{r}}\lesssim e^{-ct}\||\nabla|^{s_{1}+d}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{r}-1}_{r}},$
for all $1<r<\infty$ and $d>n/r$. This completes the proof. ∎
@@
Let $1\leq q\leq\infty,\,\,\gamma\geq 0,\,j\in\mathbb{N}$. Then, we have the following estimates for all $\varphi\in L^{1}$ and $t\gg 1$:
$\|\partial_{t}^{j}|\nabla|^{\gamma}\mathcal{G}(t,\cdot)\|_{L^{q}}\lesssim t^{-\frac{n}{2}(1-\frac{1}{q})-\frac{\gamma}{2}-j}$ (22)
and if $\gamma\in\mathbb{N}$, we have the following relation:
@@
Proof of Theorem 1.2. Firstly, we recall that the global solution to (1) satisfies the following integral equation:
$u(t,x)$ $=u^{\rm lin}(t,x)+u^{\rm non}(t,x)$
$=\varepsilon(\mathcal{K}(t,x)+\partial_{t}\mathcal{K}(t,x))\ast_{x}u_{0}(x)+\varepsilon\mathcal{K}(t,x)\ast_{x}u_{1}(x)+\int_{0}^{t}\mathcal{K}(t-\tau,x)\ast_{x}\mathcal{N}(u(\tau,x))d\tau.$
We shall use the notation and results established in Section 2 and fix $(\theta,s)$ as one of the two pairs $(\alpha,0)$ and $(2,2)$. From Lemmas 3.1 and 3.2, we can easily arrive at
@@
$\quad\quad+\varepsilon\left\||\nabla|^{s}\left(\mathcal{G}(t,x)\ast_{x}(u_{0}+u_{1})(x)-\left(\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))dx\right)\mathcal{G}(t,x)\right)\right\|_{L^{\theta}}$
$\quad\quad+\varepsilon\|\partial_{t}|\nabla|^{s}\mathcal{K}(t,x)\ast_{x}u_{0}(x)\|_{L^{\theta}}$
$\quad=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1,$
where we used Lemma 3.1 for the first term, the relation (23) for the second term, and Lemma 2.1 for the third term. Therefore, we need to prove the following relation:
$\left\||\nabla|^{s}\left(u^{\rm non}(t,x)-\left(\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\mathcal{N}(u(t,x))dxdt\right)\mathcal{G}(t,x)\right)\right\|_{L^{\theta}}=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1,$ (24)
for $(\theta,s)\in\{(2,2),(\alpha,0)\}$. Let us now divide the left-hand side term of (24) in the $L^{\theta}$ norm for $s=0$ into five parts as follows:
@@
$\quad=\int_{0}^{t/2}(\mathcal{K}(t-\tau,x)-\mathcal{G}(t-\tau,x))\ast_{x}\mathcal{N}(u(\tau,x))d\tau+\int_{t/2}^{t}\mathcal{K}(t-\tau,x)\ast_{x}\mathcal{N}(u(\tau,x))d\tau$
@@
$\lesssim\varepsilon_{0}^{1+\frac{2}{n}}(1+t)^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\int_{0}^{t/2}(1+\tau)^{-1}\mu\big(\varepsilon_{0}(1+\tau)^{-1}\big)d\tau$
@@
for $(\theta,s)=(\alpha,0)$ or $(2,2)$. Now, we consider the term $K_{2}(t,x)$. Specicially, employing again Lemma 2.3 and Lemma 2.1 for $\rho=q=\theta$ and $(s_{1},s_{2})=(s,s/2)$ we get
$\||\nabla|^{s}K_{2}(t,\cdot)\|_{L^{\theta}}$ $\lesssim\int_{t/2}^{t}(1+t-\tau)^{-\frac{s}{4}}\|\chi_{\rm L}(|\nabla|)\mathcal{N}(u(\tau,\cdot))\|_{\dot{H}^{s/2}_{\theta}}$
@@
$\lesssim\varepsilon_{0}^{1+\frac{2}{n}}\int_{t/2}^{t}(1+t-\tau)^{-\frac{s}{4}}(1+\tau)^{-\frac{n}{2}(1+\frac{2}{n}-\frac{1}{\theta})-\frac{s}{4}}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\lesssim\varepsilon_{0}^{1+\frac{2}{n}}(1+t)^{1-\frac{n}{2}(1+\frac{2}{n}-\frac{1}{\theta})-\frac{s}{2}}\mu(\varepsilon_{0}(1+t/2)^{-\frac{n}{2}})$
$=\varepsilon_{0}^{1+\frac{2}{n}}(1+t)^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}\mu(\varepsilon_{0}(1+t/2)^{-\frac{n}{2}})=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1,$ (27)
for $(\theta,s)=(\alpha,0)$ or $(2,2)$, where we note that $\mu(0)=0$, $\mu$ is continuity and $\beta_{\alpha}<1,\,\beta_{2}=0$. Taking account of $K_{3}(t,x)$ we use the mean value theorem on $t$ to get the following representation:
@@
with $\lambda_{1}=\lambda_{1}(t,\tau)\in[0,1]$. Thus, using the relation $t-\lambda_{1}\tau\sim t$ for $\tau\in[0,t/2]$ and the estimate (22), we arrive at
$\||\nabla|^{s}K_{3}(t,\cdot)\|_{L^{\theta}}$ $\lesssim\int_{0}^{t/2}\tau(t-\lambda_{1}\tau)^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\|\mathcal{N}(u(\tau,\cdot))\|_{L^{1}}d\tau$
$\lesssim\varepsilon_{0}^{1+\frac{2}{n}}t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\int_{0}^{t/2}\tau(1+\tau)^{-1}\mu\big(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\big)d\tau$
$\lesssim\varepsilon_{0}^{1+\frac{2}{n}}t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\int_{0}^{t/2}\mu\big(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\big)d\tau=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1,$ (28)
@@
$\||\nabla|^{s}K_{41}(t,\cdot)\|_{L^{\theta}}$ $\lesssim\int_{0}^{1}\int_{0}^{t/2}\int_{|y|\leq t^{1/4}}|y|\||\nabla|^{s+1}\mathcal{G}(t,x-\eta y)\|_{L^{\theta}}|\mathcal{N}(u(\tau,y))|dyd\eta$
@@
Next, using a change of variable $\lambda:=\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}$, it is a fact that
@@
Therefore, the term $K_{42}(t,x)$ will be estimated as follows:
@@
$\lesssim t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}\int_{0}^{t/2}\int_{|y|\geq t^{1/4}}\mathcal{N}(u(\tau,y))dyd\tau$
$=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1.$
In summary, we obtain
@@
4. Sharp lifespan estimate for blow-up solutions
In this section, our main aim is to prove Theorem 1.3. Following the argument in the proof of Theorem 1.1, one obtains the existence of local (in time) solutions to the Cauchy problem (1). Moreover, these solutions blow-up in finite time when the non-Dini condition (6) occurs. The remaining open problem is to determine a sharp lifespan estimate for such local solutions. Specifically, we divide the proof of Theorem 1.3 into two parts as follows.
4.1. Upper bound estimate of lifespan
In this subsection, we establish the upper bound estimate of the lifespan
$T_{\varepsilon}\lesssim\Psi^{-1}\left(C\varepsilon^{-\frac{2}{n}}\right),$
under the conditions stated in Theorem 1.3. Our proof is based on the estimates established in the proof of [8, Theorem 5]. Let us recall the test functions constructed as in [8, Theorem 5]:

### v2

@@ 2.1. Linear estimates and solution spaces
$\left|\mathcal{N}(u(\tau,\cdot))-\mathcal{N}(v(\tau,\cdot))\right|$
$\quad=\left|(u-v)(\tau,\cdot)\times\left(\int_{0}^{1}(\partial\mathcal{N})(v+\kappa(u-v))d\kappa\right)(\tau,\cdot)\right|$
$\quad\lesssim|(u-v)(\tau,\cdot)|\times\left|\left(\int_{0}^{1}|v+\kappa(u-v)|^{\frac{2}{n}}\mu\left(|v+\kappa(u-v)|\right)d\kappa\right)(\tau,\cdot)\right|$
$\quad\lesssim|(u-v)(\tau,\cdot)|\left(|u(\tau,\cdot)|^{\frac{2}{n}}+|v(\tau,\cdot)|^{\frac{2}{n}}\right)\int_{0}^{1}\|\mu(v+\kappa(u-v))(\tau,\cdot)\|_{L^{\infty}}d\kappa.$
@@
$\mathcal{J}_{1}(t)$ $:=\int_{0}^{t/2}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}(1+\tau)^{-1+\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\lesssim(1+t)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\int_{0}^{t/2}(1+\tau)^{-1+\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau\lesssim\mathcal{I}(\varepsilon_{0}),$
$\mathcal{J}_{2}(t)$ $:=\int_{t/2}^{t}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}(1+\tau)^{-1+\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\lesssim\int_{t/2}^{t}(1+t-\tau)^{-1}\mu\left(\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}\right)d\tau\lesssim\mathcal{I}(\varepsilon_{0}),$
for all $t>0$ due to $n\left(1/m-1/2\right)<2$ and $1+\tau\geq 1+t-\tau$ for $\tau\in[t/2,t]$. Thus, we complete the proof of Proposition 2.3. ∎
Proof of Theorem 1.1. Using again Lemma 2.1 and the relation (14) for the norm $L^{\infty}$, one derives
@@
for all $T>0$. For this reason, combining with Propositions 2.2 and 2.3, we obtain the following estimates for all $u,v\in X(T,\varepsilon_{0}),\,\varepsilon_{0}\in(0,1]$:
@@
for all $\varepsilon\in(0,\bar{\varepsilon}]$. Moreover, using (21) implies that $\{u_{j}\}_{j=0}^{\infty}$ is a Cauchy sequence in the Banach space $Y(T)$. Therefore, from Lemma 2.2, there exists a solution $u=\Phi[u]$ in $X(T,\varepsilon_{0})$ for all $T>0$. Since $T$ is arbitrary, the solution is global, that is, $u\in X(\infty,\varepsilon_{0})$.
Next, we need to show that $u\in\mathcal{C}([0,\infty),H^{2}\cap L^{\alpha}\cap L^{\infty})$. To prove this property, we recall the solution formula (9). Since the linear part $u^{\rm lin}(t,x)$ obviously satisfies continuity, it suffices to show that
@@
Finally, we establish the uniqueness of the global solution $u$, which lies in the space $\mathcal{C}([0,\infty),H^{2}\cap L^{\alpha}\cap L^{\infty})$. Let $u,v$ be solutions to (1) belonging to this space. For the arbitrary positive number $T$, we can see that there exists a constant $M(T)$ such that $\|u\|_{X(T)}^{2/n}+\|v\|_{X(T)}^{2/n}\leq M(T)$. Performing the same proof steps of Proposition 2.3, we obtain
@@ 3. Asymptotic behaviors of global solution
Let $1\leq\rho\leq q<\infty$, $q\neq 1$, $\beta_{q}:=(n-1)\big|\frac{1}{2}-\frac{1}{q}\big|$ and $s_{1}\geq s_{2}\geq 0$. Then, the following estimate holds for $t\geq 1$:
@@
$\hskip 28.45274pt\lesssim(1+t)^{-\frac{n}{2}(\frac{1}{\rho}-\frac{1}{q})-\frac{s_{1}-s_{2}}{2}-1}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}}+e^{-ct}\||\nabla|^{s_{1}}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{q}-1}_{q}},$
@@
$\hskip 28.45274pt\lesssim(1+t)^{-\frac{n}{2\rho}-\frac{s_{1}-s_{2}}{2}-1}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}}+e^{-ct}\||\nabla|^{s_{1}+d}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{r}-1}_{r}}.$
@@
The asymptotic behaviors in the norm $L^{q}$ with $1<q<\infty$ are obtained from Theorem 1.2 in [15]. Therefore, we need to prove the estimate of the $L^{\infty}$ norm. Proposition 2.10 in [15] implies
$\||\nabla|^{s_{1}}\big(\chi_{\rm L}(|\nabla|)\mathcal{K}(t,x)-\mathcal{G}(t,x)\big)\ast_{x}\varphi(x)\|_{L^{\infty}}$
$\quad\quad\lesssim(1+t)^{-\frac{n}{2\rho}-\frac{s_{1}-s_{2}}{2}-1}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}},$
for all $1\leq\rho\leq\infty$ and $s_{1}\geq s_{2}\geq 0$. Combining with the estimate (10), we can conclude the asymptotic behaviors in the norm $L^{\infty}$. ∎
@@
Let $1\leq q\leq\infty,\,\,\gamma\geq 0,\,j\in\mathbb{N}$. Then, we have the following estimates for all $t>0$:
$\|\partial_{t}^{j}|\nabla|^{\gamma}\mathcal{G}(t,\cdot)\|_{L^{q}}\lesssim t^{-\frac{n}{2}(1-\frac{1}{q})-\frac{\gamma}{2}-j}.$ (23)
Moreover, the following relation holds for $\gamma\in\mathbb{N}$ and $\varphi\in L^{1},\,t\gg 1$:
@@
Proof of Theorem 1.2. To begin with, we recall that the global solution to (1) satisfies the following integral equation (9). Next, we shall use the notation and results established in Section 2 and fix $(\theta,s)$ as one of the two pairs $(\alpha,0)$ and $(2,2)$. Lemmas 3.1 and 3.2 yield
@@
$\quad\qquad\quad+\varepsilon\left\||\nabla|^{s}\left(\mathcal{G}(t,x)\ast_{x}(u_{0}+u_{1})(x)-\left(\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))dx\right)\mathcal{G}(t,x)\right)\right\|_{L^{\theta}}$
$\quad\qquad\quad+\varepsilon\|\partial_{t}|\nabla|^{s}\mathcal{K}(t,x)\ast_{x}u_{0}(x)\|_{L^{\theta}}$
$\quad=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1.$
Here we used Lemma 3.1 for the first term, the relation (24) for the second term, and Lemma 2.1 for the third term. Therefore, we need to prove the following relation:
$\left\||\nabla|^{s}\left(u^{\rm non}(t,x)-\left(\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\mathcal{N}(u(t,x))dxdt\right)\mathcal{G}(t,x)\right)\right\|_{L^{\theta}}$
$\hskip 113.81102pt=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),$ (25)
for $(\theta,s)\in\{(2,2),(\alpha,0)\}$ and $t\gg 1$. Let us now divide the left-hand side term of (25) in the norm into five parts as follows:
@@
$\quad=\int_{0}^{t/2}(\mathcal{K}(t-\tau,x)-\mathcal{G}(t-\tau,x))\ast_{x}\mathcal{N}(u(\tau,x))d\tau$
$\quad\quad+\int_{t/2}^{t}\mathcal{K}(t-\tau,x)\ast_{x}\mathcal{N}(u(\tau,x))d\tau$
@@
$\lesssim\varepsilon_{0}^{1+\frac{2}{n}}(1+t)^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\int_{0}^{t/2}(1+\tau)^{-1}\mu\big(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\big)d\tau$
@@
for $(\theta,s)=(\alpha,0)$ or $(2,2)$ and $t\gg 1$. Now, we consider the term $K_{2}(t,x)$. Specicially, employing again Lemma 2.3 and Lemma 2.1 for $\rho=q=\theta$ and $(s_{1},s_{2})=(s,s/2)$ we get
$\||\nabla|^{s}K_{2}(t,\cdot)\|_{L^{\theta}}$
$\quad\lesssim\int_{t/2}^{t}(1+t-\tau)^{-\frac{s}{4}}\|\chi_{\rm L}(|\nabla|)\mathcal{N}(u(\tau,\cdot))\|_{\dot{H}^{s/2}_{\theta}}$
@@
$\quad\lesssim\varepsilon_{0}^{1+\frac{2}{n}}\int_{t/2}^{t}(1+t-\tau)^{-\frac{s}{4}}(1+\tau)^{-\frac{n}{2}(1+\frac{2}{n}-\frac{1}{\theta})-\frac{s}{4}}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\quad\lesssim\varepsilon_{0}^{1+\frac{2}{n}}(1+t)^{1-\frac{n}{2}(1+\frac{2}{n}-\frac{1}{\theta})-\frac{s}{2}}\mu(\varepsilon_{0}(1+t/2)^{-\frac{n}{2}})$
$\quad=\varepsilon_{0}^{1+\frac{2}{n}}(1+t)^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}\mu(\varepsilon_{0}(1+t/2)^{-\frac{n}{2}})=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1,$ (28)
where we note that $\mu(0)=0$, $\mu$ is continuity and $\beta_{\alpha}<1,\,\beta_{2}=0$. Taking account of $K_{3}(t,x)$ we use the mean value theorem on $t$ to get the following representation:
@@
with $\lambda_{1}=\lambda_{1}(t,\tau)\in[0,1]$. Thus, using the relation $t-\lambda_{1}\tau\sim t$ for $\tau\in[0,t/2]$ and the estimate (23), the following estimates hold for $t\gg 1$:
$\||\nabla|^{s}K_{3}(t,\cdot)\|_{L^{\theta}}$
$\quad\lesssim\int_{0}^{t/2}\tau(t-\lambda_{1}\tau)^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\|\mathcal{N}(u(\tau,\cdot))\|_{L^{1}}d\tau$
$\quad\lesssim\varepsilon_{0}^{1+\frac{2}{n}}t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\int_{0}^{t/2}\tau(1+\tau)^{-1}\mu\big(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\big)d\tau$
$\quad\lesssim\varepsilon_{0}^{1+\frac{2}{n}}t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}-1}\int_{0}^{t/2}\mu\big(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\big)d\tau=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),$ (29)
@@
$\||\nabla|^{s}K_{41}(t,\cdot)\|_{L^{\theta}}$ $\lesssim\int_{0}^{1}\int_{0}^{t/2}\int_{|y|\leq t^{1/4}}|y|\||\nabla|^{s+1}\mathcal{G}(t,x-\eta y)\|_{L^{\theta}}|\mathcal{N}(u(\tau,y))|dyd\tau d\eta$
@@
Next, employing a change of variable $\lambda:=\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}$, we gain
@@
Therefore, thanks to (23), the term $K_{42}(t,x)$ will be estimated as follows:
@@
$\lesssim t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}\int_{0}^{t/2}\int_{|y|\geq t^{1/4}}\mathcal{N}(u(\tau,y))dyd\tau=o(t^{-\frac{n}{2}(1-\frac{1}{\theta})-\frac{s}{2}}),\quad t\gg 1.$
In summary, we gain
@@
4. Some remarks to sharp lifespan estimates
Following the argument in the proof of Theorem 1.1, one obtains the existence of global (in time) solutions to the Cauchy problem (1). On the other hand, Theorem 5 in [9] shows that the solution to (1) blows up in finite time for all $n\geq 1$ when the non-Dini condition (6) occurs. Now, we define the following function:
Proposition 4.1 (Sharp lifespan estimates).